The sum of three times a first number and twice a second number is 8. If the second number is subtracted from twice the first number, the result if 3. Find the numbers.
Answer by mananth(16949)

first number = x
second number = y
...
3x+2y=8.......................1
2x-y=3........................2
multiply equation2 by 2 and add it to equation 1
4x-2y=6
add to 1
7x=14
/7
x= 2
...
plug the value of x in equation 2
2*2-y=3
4-y=3
y=1
...
m.ananth@hotmail.ca